### v1

# See What You Seek: Semantic Contextual Integration for Cloth-Changing Person Re-Identification

Abstract
Cloth-changing person re-identification (CC-ReID) aims to match individuals across multiple surveillance cameras despite variations in clothing. Existing methods typically focus on mitigating the effects of clothing changes or enhancing ID-relevant features but often struggle to capture complex semantic information. In this paper, we propose a novel prompt learning framework, Semantic Contextual Integration (SCI), for CC-ReID, which leverages the visual-text representation capabilities of CLIP to minimize the impact of clothing changes and enhance ID-relevant features. Specifically, we introduce Semantic Separation Enhancement (SSE) module, which uses dual learnable text tokens to separately capture confounding and clothing-related semantic information, effectively isolating ID-relevant features from distracting clothing semantics. Additionally, we develop a Semantic-Guided Interaction Module (SIM) that uses orthogonalized text features to guide visual representations, sharpening the model’s focus on distinctive ID characteristics. This integration enhances the model’s discriminative power and enriches the visual context with high-dimensional semantic insights. Extensive experiments on three CC-ReID datasets demonstrate that our method outperforms state-of-the-art techniques. The code will be released at github.
Index Terms:
Person Re-Identification, Clothing Changes, Vision-Language Models, Prompt Learning, Semantic Integration.I Introduction
Person re-identification (Re-ID) is a critical task that involves identifying individuals across different locations over time, with significant applications in video surveillance and smart city infrastructure [1, 2, 3]. Traditional Re-ID methods [4, 5, 6, 7] focus on clothing features, such as texture and color, to improve performance. However, these methods become ineffective when individuals change clothes, highlighting the need for robust approaches that can handle clothing variations in real-world settings.
To address the challenge of clothing changes in Re-ID (CC-ReID), extensive research has yielded promising results [8, 9]. Existing methods can be broadly categorized into two categories: those aiming to reduce the influence of clothing and those focusing on enhancing ID-relevant information. The first category seeks to minimize the impact of clothing features, as illustrated in Fig. 1(a). For instance, CAL [10] uses adversarial learning to reduce reliance on clothing by penalizing predictions based on the RGB modality, while AIM [11] employs causal intervention with a dual-branch model to mitigate clothing bias. The second category emphasizes ID-relevant features using auxiliary cues such as pose, gait, or human parsing networks. For example, GI-ReID [12] uses gait information to learn clothing-agnostic representations, and SCNet [13] leverages human body parsing network to learn identity-relevant features through semantic consistency constraints.
Although these methods address clothing changes, they typically focus either on minimizing clothing influence or enhancing ID cues through explicit features like body contours. We argue that implicit positive factors such as hair, glasses, or backpacks are embedded in the visual features but are challenging to define explicitly or separately. As shown in Fig. 1(b), our goal is to remove negative influences while making implicit positive factors more explicit.
Recent advancements in visual-language learning have shown great potential in enhancing visual recognition and understanding. Contrastive Language-Image Pre-training (CLIP) [14], a prominent cross-modal pre-training model, has effectively bridged the gap between visual data and natural language, enabling more context-aware and interpretatively rich analysis. CLIP has achieved remarkable success in various downstream visual tasks [15, 16]. In Re-ID, recent work [17, 18] demonstrates that linking visual content with corresponding language descriptions allows models to capture complex semantics associated with individuals. However, these studies have primarily focused on traditional Re-ID, applying CLIP to CC-ReID has not yet been explored. This motivates us to leverage prompt learning to reveal high-dimensional semantic information, thereby enhancing visual-textual connections and making implicit positive factors more explicit.
In this paper, we propose the Semantic Contextual Integration (SCI) framework for CC-ReID, which utilizes CLIP to capture high-dimensional semantic features. Since CLIP struggles to selectively ignore aspects like clothing through negative prompts, we introduce a Semantic Separation Enhancement (SSE) module to minimize the influence of clothing by isolating semantic-level negative factors and preserving key positive features for CC-ReID.
Specifically, we use dual learnable text tokens to separately capture confounding and clothing-related semantic information. We then apply an orthogonalization process on two distinct sets of text features to filter out clothing-related features from the confounding ones. SSE not only reduces the negative impact of clothing but also retains important positive semantic factors. These refined text features are used to guide the visual encoder.
To further emphasize implicit positive factors in visual images, we introduce a Semantic-Guided Interaction Module (SIM), which directs the visual encoder’s extraction process using the filtered text features. This allows the model to focus on ID-relevant elements beyond explicit features such as body contours. By leveraging CLIP’s text-visual integration capabilities, our method enhances the interaction between the two modalities, improving visual feature refinement and multi-modal alignment.
Our contributions are summarized fourfold:
We propose the Semantic Contextual Integration (SCI) framework to leverage semantic information in CC-ReID, removing negative factors, emphasizing implicit positive elements, and refining visual representations.
We introduce the Semantic Separation Enhancement (SSE) module to filter and refine text-level features, improving the model’s ability to isolate key semantic information.
We design the Semantic-Guided Interaction Module (SIM) to guide visual representations using refined text features, enhancing multi-modal integration and alignment.
Extensive experiments on three standard CC-ReID datasets demonstrate the effectiveness of our approach, providing a robust solution to real-world Re-ID challenges.
II Related Work
II-A Cloth-Consistent Person Re-ID
Person Re-ID under consistent clothing conditions has been extensively studied [19, 20, 21]. These methods primarily rely on clothing color and texture to extract discriminative features. Feature representation learning in this context can be categorized into three types: global features, local features, and auxiliary information-based features.
Global feature representation methods extract a single feature vector for each person image [22]. Local feature representation methods aggregate part-based features to address misalignments, using techniques like human parsing [23] for semantic part detection, horizontal division [24] for structured feature aggregation, hard region-level and soft pixel-level attention [25] for optimized features, and IANet [26] for enhanced small visual cues. Auxiliary information-based methods incorporate additional data, such as semantic attributes [19] or generated/augmented samples [20], to enhance generalization by embedding richer contextual information. And ISP [27] locates human body parts and personal belongings at pixel-level for aligned with labels.
Although effective in handling variations in appearance, background, and camera perspectives, these methods assume consistent clothing, leading to performance declines in scenarios with clothing changes, which are common in real-world applications. This limitation has spurred increased focus on addressing cloth-changing person Re-ID challenges.
II-B Cloth-Changing Person Re-ID
To advance CC-ReID, several datasets such as PRCC [28], LTCC [29], Celeb-reID [30], and VC-Clothes [31] have been introduced, providing diverse data for effective model evaluation. The core challenge in CC-ReID is to learn features that reliably identify individuals despite clothing variations.
There are three primary approaches to this challenge. The first leverages auxiliary soft-biometric information. For example, CAMC [32] incorporates body shape semantics into ID-related features, enhancing robustness to clothing changes, while FSAM [33] learns fine-grained body shape features to complement clothing-independent features. M2Net [34] utilizes contour and human parsing information to learn features that are robust to appearance changes.
The second approach involves disentangling features to isolate clothing-irrelevant information and reduce clothing bias. For instance, CAL [10] uses clothes-based adversarial loss to emphasize clothing-irrelevant features, and AIM [11] applies causal intervention to automatically eliminate clothing effects.
The third approach employs data augmentation or adaptation techniques. Pos-Neg [8] and CCFA [9] enhance clothing color and texture diversity without additional data collection, improving robustness to clothing variations. RCSANet [35] explicitly constructs a clothing status, which is used to enhance the robustness of ID features for handling both cloth-changing and no-cloth-changing cases.
Despite these advancements, existing methods primarily rely on visual inputs for model training, limiting their capacity for deep semantic understanding. In this paper, we propose a novel method that integrates visual and natural language information, utilizing CLIP to enhance the interaction of the visual content with corresponding textual descriptions.
II-C Vision-Language Learning
Vision-language pre-training (VLP) has recently shown significant promise in enhancing the performance of various downstream vision tasks by training models to associate images with textual information. CLIP [14], a leading VLP model, employs paired image and text encoders to bridge the gap between natural language and visual content. By using similarity learning, CLIP aligns textual and visual representations in a shared space, enhancing tasks such as image captioning and classification.
Building on CLIP, recent advances in prompt learning, such as CoOp [36], transform prompt context words into learnable vectors. CoCoOp [37] extends this by generating input-conditional tokens for each image through a lightweight network. DenseCLIP [38] adapts this approach for dense prediction models, introducing pixel-text matching to utilize contextual information from a pre-trained language model.
In person Re-ID, CLIP-ReID [17] is the first to use ID-specific learnable tokens and a two-stage training strategy to fully utilize text for specifying visual concepts. CCLNet [18] advances this approach with a novel prompt learning paradigm for unsupervised visible-infrared Re-ID. However, these studies primarily focus on scenarios with consistent clothing, leaving CLIP’s potential in CC-ReID largely unexplored. In this paper, we propose the CLIP-CCReID framework to mitigate the impact of clothing changes in CC-ReID.
III Proposed Method
III-A Preliminaries and Overview
The large-scale pre-trained vision-language model CLIP [14] demonstrates significant advantages in learning joint representations of images and text. CLIP consists of two encoders: a visual encoder and a text encoder . CLIP aligns the embedding spaces of visual and textual data during pre-training using a contrastive objective.
For downstream classification tasks, a simple and effective method [36] constructs text prompts using a template like “a photo of a [CLS].”, where [CLS] is replaced with specific class names. Given an image, CLIP computes similarities between the image embedding and the text prompt embeddings, the process involves transforming the image and text into a shared high-dimensional space where their relationships can be quantitatively assessed. CLIP then selects the class corresponding to the highest similarity as the prediction.
Building upon CLIP, CoOp [36] introduces learnable textual contexts, represented as “ [CLS].”, where () are learnable tokens. This approach enhances the model’s transferability and performance. Leveraging the capabilities of CLIP and CoOp, we employ these textual descriptions to achieve deeper semantic understanding for CC-ReID, as illustrated in Fig. 2, enabling more robust and accurate modeling despite clothing changes.
III-B Semantic Separation Enhancement
Since CLIP cannot learn negative semantics or ignore specific elements like clothing through negative prompts, we propose the Semantic Separation Enhancement (SSE) module to address this limitation. This module mitigates the impact of clothing variations by subtracting negative semantic elements, preserving only the high-dimensional positive features crucial for CC-ReID. The complete algorithm is summarized in Algorithm 1.
We define dual textual prompts as follows:
where () are learnable tokens initialized randomly, and denotes the number of tokens. In the first training stage, we use the pre-trained visual encoder and text encoder to extract image and dual text features, freezing the encoders’ parameters to focus on optimizing the text tokens . This allows us to learn contextual representations by updating the text tokens, thereby acquiring distinct textual representations for each ID and their clothing:
where and represent the ID-specific text features and clothing text features, respectively.
After extracting the dual text representations, we project onto to reduce the influence of clothing and emphasize the implicit positive factors. This step clarifies the relation between clothing representations and ID representations, thereby minimizing the impact of clothing variations. The computation process is as follows:
We then perform orthogonalization by subtracting the projection from , removing the component in the direction of at the textual level. This process ensures that the features are aligned with the positive features while removing the influence of negative factors. The orthogonalization process is:
where is the feature orthogonalized to align with and orthogonal to .
To optimize the text features, we design a loss function that maximizes the similarity between and , while minimizing the similarity with . This enables the model to better capture positive factors without being affected by negative aspects, such as clothing changes. The similarity calculations are:
where denotes the average value, and represents the cosine similarity function. The loss function is as follows:
where and are weighting factors controlling the contributions of ID and clothing similarities.
Finally, we calculate the image-to-text contrastive loss :
where and are paired visual and text embeddings, denotes the similarity function, and is the batch size. Since multiple images in a batch may belong to the same ID, meaning there may be multiple positive samples, the text-to-image contrastive loss is calculated as:
where is the set of all positive indices for in the batch, and is the text embedding corresponding to label . Therefore, the overall loss function is:
III-C Semantic-Guide Interaction Module
Previous work [10, 17, 39] shows that reducing clothing bias in visual or textual features improves performance in cloth-changing scenarios. However, these methods typically treat visual and textual branches independently, lacking interaction. Our method refines visual representations by leveraging clothing-irrelevant textual features, enhancing the interaction between visual and textual branches. This integration introduces more robust and invariant descriptors into the visual processing, improving the model’s ability to identify individuals across clothing changes. The algorithm is summarized in Algorithm 2.
As illustrated in Fig. 3, we employ a cross-attention mechanism in a Transformer decoder [40] to model interactions between visual and textual data. Instead of traditional self-attention, we use a non-local operation [41] to capture global context, ensuring the model can integrate information from distant but relevant positive parts of the visual image, which is crucial in cloth-changing scenarios. During this stage, we update the visual encoder while keeping other components frozen. The computation is as follows:
where is the input feature map, , , and are linear transformations, and is the number of elements in the feature map. captures global context information.
The final visual feature map is:
where contains learnable parameters. is the final output feature map obtained by adding this context-aware feature to the original feature map.
We further enhance interaction by using text embeddings as queries and visual embeddings as keys and values. This cross-attention mechanism refines visual features using text information, ensuring alignment with relevant semantic context:
where is the dimensionality of the keys. This implementation strategically encourages the text features to find the most relevant visual clues.
Finally, we update the visual features. This process ensures that the visual features are enriched with relevant semantic information from the text, leading to more robust and contextually aware visual representations:
where is a learnable parameter controlling the degree of refinement applied to the visual features.
This module employs an interaction-to-alignment approach, ensuring that text embeddings guide the visual encoding process, transforming implicit information into explicit visual cues. By emphasizing positive factors and reducing negative ones, this method enhances the model’s ability to consistently identify individuals across various clothing scenarios.
For the objective function, we incorporate cross-entropy loss and clothes-based adversarial loss [10] to optimize the visual encoder:
where is the true label for the -th sample, and is the predicted probability of the true label . The clothes-based adversarial loss is expressed as:
where is the number of clothing categories, is the clothes classifier, is the weight for the -th class, is the feature of sample , and is a temperature parameter.
To fully leverage the capabilities of CLIP, we calculate the image-to-text cross-entropy loss as:
where label smoothing is applied to . The total loss for the SIM module is:
IV Experimental Results
IV-A Datasets and Evaluation Metrics
We evaluate our proposed method on three standard cloth-changing datasets: LTCC [29], PRCC [28], and VC-Clothes [31], following CAL [10] and AIM [11]. Additionally, we also evaluate our method on two traditional Re-ID datasets: Market-1501 [42], and MSMT17 [43], to validate its applicability.
LTCC dataset contains 17,138 images of 152 IDs captured by 12 cameras. The training set comprises 9,576 images of 77 IDs, and the testing set includes 7,543 images (493 queries and 7,050 gallery images) of 75 IDs. This long-term Re-ID dataset features frequent clothing changes and varied environmental conditions, with each ID having between 2 and 14 outfits.
PRCC dataset consists of 33,698 images of 221 IDs captured by three cameras. The training set contains 22,898 images of 150 IDs, while the testing set includes 10,800 images of 71 IDs. Each person has two outfits, with cameras A and B capturing the same clothes and camera C capturing a different outfit.
VC-Clothes dataset is a synthetic dataset generated using the GTA5 game engine. It comprises 9,449 images of 256 IDs across both training and testing sets, featuring 1,241 clothing items across four camera views.
Market-1501 dataset includes 1,501 individuals recorded by six different cameras. It consists of 12,936 images of 751 IDs for training and 19,732 images of 750 IDs for testing. The evaluation is conducted using the single query setting.
MSMT17 dataset includes 1,041 individuals captured by 15 different cameras. The training set contains 32,621 images of 1,041 IDs, while the testing set includes 51,027 images of 11,659 IDs. This dataset is notably large and complex.
We use the cumulative matching characteristic (CMC) curve and mean average precision (mAP) as evaluation metrics. Rank- in CMC measures the likelihood of retrieving the correct ID in the top- results, while mAP captures the average retrieval performance across all queries.
Evaluation is conducted under three settings: 1) General setting: Excludes same-ID and same-camera samples, using both cloth-changing and clothing-consistent samples. 2) Same-clothes setting: Excludes same-ID and same-camera samples, using only clothing-consistent samples. 3) Cloth-changing setting: Excludes same-ID, same-camera, and same-clothing samples, using only cloth-changing samples.
IV-B Implementation Details
We adopt a modified ResNet-50 pre-trained on CLIP as the backbone for feature extraction. Images are resized to pixels, with a batch size of 64. Training employs the Adam optimizer [44] with random horizontal flipping, cropping, and erasing [45] for data augmentation. A global attention pooling layer reduces the feature dimension from 2048 to 1024, matching the text feature dimension, which is scaled from 512 to 1024.
In the first stage, two prompt learners are trained for 60 epochs on LTCC, VC-Clothes, Market-1501, and MSMT17, and 20 epochs on PRCC, with an initial learning rate of adjusted using a cosine decay schedule. In the second stage, the visual encoder is trained for 120 epochs on LTCC, VC-Clothes, Market-1501, and MSMT17, and 20 epochs on PRCC, with learning rates initially set to and reduced by a factor of 10 at the 40th and 70th epochs. All experiments are conducted on the PyTorch platform using a single A100 GPU.
IV-C Comparison with State-of-the-Art Methods
We compare our method with state-of-the-art techniques on LTCC, PRCC, and VC-Clothes, including traditional Re-ID methods HACNN [25], PCB [24], IANet [26], and ISP [27], and specialized CC-ReID methods FSAM [33], RCSANet [35], CAL [10], GI-ReID [12], M2Net [34], and AIM [11].
| Methods | Venue | LTCC | PRCC | VC-Clothes | |||||||||
| General | Cloth-changing | Same-clothes | Cloth-changing | General | Cloth-changing | ||||||||
| R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | ||
| HACNN [25] | CVPR ’18 | 60.2 | 26.7 | 21.6 | 9.3 | 82.5 | 84.8 | 21.8 | 23.2 | 68.6 | 69.7 | 49.6 | 50.1 |
| PCB [24] | ECCV ’18 | 65.1 | 30.6 | 23.5 | 10.0 | 99.8 | 97.0 | 41.8 | 38.7 | 87.7 | 74.6 | 62.0 | 62.2 |
| IANet [26] | CVPR ’19 | 63.7 | 31.0 | 25.0 | 12.6 | 99.4 | 98.3 | 46.3 | 45.9 | - | - | - | - |
| ISP [27] | ECCV’20 | 66.3 | 29.6 | 27.8 | 11.9 | 92.8 | - | 36.6 | - | 94.5 | 94.7 | 72.0 | 72.1 |
| FSAM [33] | CVPR’21 | 73.2 | 35.4 | 38.5 | 16.2 | - | - | - | - | 94.7 | 94.8 | 78.6 | 78.9 |
| RCSANet [35] | ICCV’21 | - | - | - | - | 100.0 | 97.2 | 50.2 | 48.6 | - | - | - | - |
| CAL [10] | CVPR’22 | 73.4 | 39.4 | 38.0 | 17.2 | 100.0 | 99.6 | 54.7 | 55.4 | 93.1 | 88.3 | 82.6 | 81.7 |
| GI-ReID [12] | CVPR’22 | 63.2 | 29.4 | 23.7 | 10.4 | - | - | - | - | - | - | 64.5 | 57.8 |
| M2Net [34] | ACM MM’22 | - | - | - | - | 99.5 | 99.1 | 59.3 | 57.7 | - | - | - | - |
| AIM [11] | CVPR’23 | 75.7 | 41.3 | 41.8 | 17.9 | 100.0 | 99.7 | 56.3 | 56.5 | 91.2 | 84.9 | 81.0 | 75.7 |
| CLIP-ReID [17] | AAAI’23 | 73.0 | 36.5 | 38.3 | 15.9 | 100.0 | 99.5 | 57.0 | 55.3 | 93.1 | 86.9 | 85.9 | 79.3 |
| SCI (Ours) | – | 75.7 | 40.6 | 42.1 | 18.6 | 99.6 | 97.7 | 59.8 | 56.2 | 94.9 | 89.2 | 89.2 | 83.1 |
As shown in Table I, CC-ReID methods significantly outperform general Re-ID methods, primarily because general methods rely heavily on clothing for identification, which is less effective in long-term surveillance with clothing changes. Our SCI method achieves competitive results by effectively eliminating ID-irrelevant features and enhancing positive ID elements in visual images.
On LTCC, SCI surpasses the baseline [17] in the cloth-changing setting, with improvements of 3.8% in Rank-1 accuracy and 2.7% in mAP. Although our mAP is slightly lower than AIM [11] in the general setting, this is likely due to LTCC’s inherent challenges, where models often rely on clothing information. By filtering out such information, SCI experiences minor declines in non-cloth-changing scenarios, but these are negligible given the overall benefits. Notably, SCI achieves these results without incorporating additional modalities or modules, unlike other methods. The primary advantage of our method lies in its ability to avoid carrying clothing-relevant information, which is often unavoidable in methods relying on auxiliary data such as poses and gaits. Compared with non-auxiliary-based methods, SCI outperforms CAL by 12.7% in Rank-1 and 6.3% in mAP in the cloth-changing setting on VC-Clothes. On PRCC, SCI also surpasses AIM by 3.5% in Rank-1 accuracy for clothing changes. Although the differences on other metrics are less pronounced due to PRCC’s simpler structure (only two outfits per ID), SCI still shows notable improvements. On VC-Clothes, SCI demonstrates significant gains in the cloth-changing setting, with a 7.4% higher mAP than AIM, underscoring the effectiveness of our approach in challenging scenarios.
IV-D Ablation Studies and Analysis
We conduct ablation studies to assess the effectiveness of the components in SCI and analyze parameter impacts.
IV-D1 Effectiveness of Semantic Separation Enhancement
We first evaluate the baseline without additional modules, where the model relies solely on a prompt “a photo of a person” for classification. This approach shows reduced performance. Introducing the SSE module, which uses two prompts (“a photo of a person” and “a photo of the clothes”) to learn features separately at the textual level, improves Rank-1 and mAP by around 2.0% across datasets, demonstrating its ability to filter clothing features.
These results indicate: 1) The original generic textual representations are confounding, containing both ID and clothing features, leading the model to focus excessively on cloth-relevant aspects. 2) Separating these features at the textual level is simple yet effective. These findings demonstrate the effectiveness of setting specific prompts for each factor and the robustness of our SSE module against both cloth-changing and general scenarios.
| Baseline | SSE | SIM | LTCC | PRCC | VC-Clothes | |||||||||
| General | Cloth-changing | Same-clothes | Cloth-changing | General | Cloth-changing | |||||||||
| R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | |||
| ● | ○ | ○ | 73.0 | 36.5 | 38.3 | 15.9 | 100.0 | 99.5 | 57.0 | 55.3 | 93.1 | 86.9 | 85.9 | 79.3 |
| ● | ● | ○ | 76.5 | 39.7 | 39.8 | 18.3 | 100.0 | 98.4 | 59.0 | 54.7 | 94.1 | 88.0 | 87.6 | 81.1 |
| ● | ○ | ● | 72.8 | 36.3 | 37.8 | 16.1 | 99.7 | 98.0 | 52.4 | 49.9 | 94.0 | 88.9 | 88.4 | 82.4 |
| ● | ● | ● | 75.7 | 40.6 | 42.1 | 18.6 | 99.6 | 97.7 | 59.8 | 56.2 | 94.9 | 89.2 | 89.2 | 83.1 |
IV-D2 Effectiveness of Semantic-Guided Interaction Module
We further assess SIM by integrating it with the baseline model. As shown in Table II, using SIM alone often degrades performance, particularly in the cloth-changing scenario on PRCC, where Rank-1 drops by 4.6% and mAP decreases from 55.3% to 49.9%.
The decline in performance in this set of ablation experiments is primarily due to the absence of the SSE module. Typically, the SSE module generates final textual features that guide the visual representations during the updating process of the visual encoder in SIM. However, with only a single prompt “a photo of a person” used to extract textual representations, many ID-irrelevant negative factors, particularly clothing factors, are included. This results in an excessive emphasis on clothing-relevant representations during guidance in SIM, reducing overall performance.
| (a) Baseline (stage1) | (b) Baseline (stage2) | (c) Ours (stage1) | (d) Ours (stage2) |
IV-D3 Feature Distribution Analysis
To illustrate SCI’s effectiveness, we employ t-SNE visualization [46], as depicted in Fig. 4. This visualization plots the statistical distribution for 20 randomly selected categories from LTCC, comparing the latent space distributions at different stages of the baseline model with our approach.
From Figs. 4(a) and (c), it is observable that both the baseline and our method exhibit disorganized feature distributions with blurred class boundaries. During this stage, only the prompts are trained while the visual and text encoders remain frozen, limiting their capability to effectively extract diverse representations. In contrast, as shown in Figs. 4(b) and (d), the feature distributions for both methods become clearer, indicating the effectiveness of the representations derived from the text-image similarity computations.
Notably, the clusters in Fig. 4(d) are more compact and distinct compared to those in Fig. 4(b). The red and green dashed circles in Fig. 4(b) enclose samples with considerable scatter, which are better clustered together in Fig. 4(d). Therefore, this visualization not only serves as compelling validation of our capability to effectively extract features and discriminate IDs but also highlights our method’s potential to significantly advance the field in cloth-changing scenarios.
IV-D4 Visualization of Feature Maps
To illustrate the distinctions between the baseline [17] and SCI, we present heatmaps of regions of interest in Fig. 5, demonstrating performance across LTCC, VC-Clothes, and PRCC. Fig. 6 shows VC-Clothes for the same ID but with different outfits, illustrating the regions of interest that our SCI model focuses on under clothing variations. These visualizations provide insights into the unique differences in feature activation patterns.
As shown in Fig. 5, the baseline method tends to focus on more dispersed areas, which include negative factors, potentially detracting from its effectiveness. Conversely, SCI emphasizes highly discriminative ID features, such as the person’s head, shoulders, lower body, and shoes. This specific emphasis on footwear aligns with findings from previous studies [47], which suggest that footwear often remains consistent across different outfits and scenarios. This indicates our method’s enhanced ability to capture crucial positive factors when encountering challenging scenarios.
In Fig. 6, for each person, the corresponding heatmaps illustrate how the feature extraction model highlights various regions of interest across different outfits, reflecting the model’s focus on certain body parts regardless of clothing changes. However, there are some limitations. As indicated by the red circles, the model tends to focus excessively on areas like the knees. If people wear long pants, this focus could potentially affect the results. We are aware of this issue and will continue to optimize our model to address it.
| (a) R-1 Accuracy on cloth-changing | (b) mAP Accuracy on cloth-changing | (c) R-1 Accuracy on general | (d) mAP Accuracy on general |
IV-D5 Comparative Analysis of Mechanisms in SIM
As shown in Fig. 7, we evaluate the effectiveness of the non-local operation [41] compared to self-attention mechanisms [40] in leveraging textual information obtained from the SSE module to guide visual representations. In cloth-changing scenarios, the non-local mechanism yields superior performance, achieving 42.1% in Rank-1 accuracy and 18.6% in mAP, as depicted in Figs. 7(a) and (b), thereby outperforming the self-attention mechanism by margins of 4.6% and 2.4%, respectively.
This underscores the non-local mechanism’s enhanced capability to capture comprehensive global context, which is vital for integrating information from disparate yet significant segments of the visual field in cloth-changing scenarios. In general scenarios, as illustrated in Figs. 7(c) and (d), both mechanisms exhibit comparable efficacy, with the non-local mechanism slightly edging out self-attention by a 2.0% increase in mAP. This suggests the non-local mechanism is robust and can be effectively deployed for person Re-ID tasks.
IV-D6 Loss Functions Analysis
We conduct experimental evaluations to explore the effectiveness of (9), as demonstrated in Fig. 8(a). This loss function significantly improves performance by optimizing for higher ID similarity and reducing clothing similarity, making it particularly effective for applications that need to distinguish personal ID from external features like clothing, relevant in tasks such as cloth-changing scenarios. It helps the model focus on positive factors while ignoring negative variations such as clothing changes, thereby enhancing adaptability and accuracy in complex environments.
| Methods | Venue | Market1501 | MSMT17 | |||
| R-1 | mAP | R-1 | mAP | |||
| Cloth- consistent | AGW [48] | TPAMI’21 | 95.1 | 87.8 | 68.3 | 49.3 |
| reID-NAS [49] | TNNLS’22 | 95.1 | 85.7 | 79.5 | 53.3 | |
| CLIP-ReID [17] | AAAI’23 | 95.7 | 89.8 | 84.4 | 63.0 | |
| IRM [50] | CVPR’24 | 96.5 | 93.5 | 86.9 | 72.4 | |
| Cloth- changing | CAL [10] | CVPR’22 | 85.0 | 67.8 | 65.1 | 38.4 |
| AIM [11] | CVPR’23 | 70.3 | 48.0 | 26.6 | 12.0 | |
| SCI (Ours) | – | 85.7 | 68.6 | 73.5 | 47.4 | |
IV-D7 Cloth-consistent Re-ID Scenario
To validate the applicability of our method in cloth-consistent scenarios, we conducted experiments using our cloth-changing approach on two traditional ReID datasets: Market1501 and MSMT17. For ease of comparison, we also incorporated cloth-consistent methods: AGW [48], reID-NAS [49], CLIP-ReID [17], and IRM [50]. This enables a comprehensive evaluation of the performance in varying scenarios.
As shown in Table III (cloth-consistent), existing Re-ID methods designed for traditional datasets have achieved progressively better performance, even on larger-scale datasets like MSMT17. While our method is specifically designed to address the more realistic scenario of clothing changes, we are also interested in evaluating its performance on conventional datasets. From Table III (cloth-changing), it can be observed that our SCI method achieves the best performance on both Market1501 and MSMT17 among cloth-changing methods. Although there is still a gap compared to cloth-consistent methods, this highlights an important direction for future research and improvement.
IV-D8 Visualization of Retrieved Examples
We present the visualization of the top-10 retrieved results from randomly selected query examples on LTCC in Fig. 9. It observes that the retrieval results are predominantly optimistic, reflecting a high level of accuracy and relevance in most cases. However, some incorrect results still persist, which could be due to challenges such as varying viewpoints, low lighting conditions, or occlusions within the images. These issues underscore potential areas for future research. Despite these challenges, the overall result strongly supports the effectiveness of our SCI, demonstrating its robustness across a variety of scenarios.
IV-D9 Parameter Selection
We evaluate the parameters and specified in (9). The results for LTCC are displayed in Fig. 8(b). We determine that and provide the best performance in most scenarios, optimizing the balance between enhancing ID and reducing clothing similarity, while feature orthogonalization minimizes interference between ID and clothing features. This setting allows the model to effectively discern between the positive representations and their negative factors, such as clothing. Additionally, and obtain the second-best results. This analysis underscores the effectiveness and adaptability of our proposed method in real-world surveillance applications.
IV-D10 Efficiency Analysis
To verify the efficiency and performance of SCI, we analyze several computational and training metrics across different methods, detailed in Table IV. AIM [11] utilizes a dual-branch ResNet model, achieving an R-1 accuracy of 41.8%, which surpasses CAL [10], another ResNet-based model. Although AIM requires higher GFLOPs compared to CAL, it significantly outperforms CAL in terms of accuracy with the same parameter count.
In CLIP-ReID [17] and our method, we integrate textual modality information, transitioning from a single-modality to a multimodal approach. This justifies the observed increases in both GFLOPs and parameters, which remain within acceptable limits. Our SCI achieves an R-1 accuracy of 42.1%, demonstrating the value of incorporating textual information. This allows the model to better distinguish between different IDs, especially in challenging scenarios. Overall, our method provides a good balance between computational cost and performance enhancements.
V Conclusion and Limitation
In this study, we propose the Semantic Contextual Integration (SCI) network, a novel method incorporates prompt learning to address cloth-changing person re-identification (CC-ReID). The SCI network includes a Semantic Separation Enhancement (SSE) module that generates ID-specific and clothing prompts to effectively isolate clothing semantics (negative factors) from ID information, thereby preserving high-dimensional semantic details (positive factors) embedded in visual images. Additionally, the Semantic-Guided Interaction Module (SIM) leverages textual cues to guide and enhance the discrimination of visual features, facilitating seamless interaction between visual and textual branches. We believe this work provides valuable insights and potential directions for future research into the role of prompt learning in advancing CC-ReID.
While learnable prompts enable the model to autonomously construct and learn, have indeed demonstrated superior outcomes and a significant enhancement in performance compared to the manually crafted prompts used in CLIP, the prompts derived by the model may lack semantic clarity and remain illegible to human observers. As a result, it is not feasible to visually represent the learned content in a textual format. Thus, future research will need to address the interpretability of these prompts and investigate methodologies that enhance their comprehensibility while maintaining their efficacy.